OpenStudy (anonymous):

A rectangular prism is truncated removing 10% of its height. The original height of the prism was 100 meters and it has a length of 4 meters and a width of 20 meters. What is the volume of the truncated prism? 800 m3 4,200 m3 3,800 m3 7,200 m3

OpenStudy (apoorvk):

 in this prism problem, you basically are just cutting off the upper shaded part from th eprism that mean s you are removing the 10 percen tof it, of the new volume v'= v - v/10 =0.9v = 0.9(8000) cubic meters = 7,200 cubic meters.